OpenStudy (anonymous):

I need some help to solve for x. 1/5(2x-10) + 4x = -3(1/5x +4)

OpenStudy (xapproachesinfinity):

is that \[\frac{1}{5}(2x-10)+4x=-3(\frac{1}{5}x+4)\]

OpenStudy (xapproachesinfinity):

is that what you wrote or perhaps something else the way you wrote has a different meaning

OpenStudy (anonymous):

Yes that is. :)

OpenStudy (xapproachesinfinity):

ok first distribute 1/5 can you do that?

OpenStudy (xapproachesinfinity):

there is no point if you are not trying at least?

OpenStudy (anonymous):

I'm not sure what distribute is. :(

OpenStudy (diamondboy):

what he is trying to say is multiply everything in d parentheses by 1/5

OpenStudy (xapproachesinfinity):

i'm you have done something like this \[a(b+c)=ab+ac\]

OpenStudy (anonymous):

Oh.. okay, one minute I have to figure it out on paper. :)

OpenStudy (anonymous):

Sorry that I took a while. 10 x 1/5 = 2. This is new to me lol

OpenStudy (diamondboy):

yep........ u are doing good @jadamarierose

OpenStudy (anonymous):

Thanks @diamondboy

OpenStudy (diamondboy):

uwc but that's not all

OpenStudy (anonymous):

I know there's more :D

OpenStudy (diamondboy):

:D....so after u distribute 1/5 what do u get?

OpenStudy (anonymous):

2

OpenStudy (diamondboy):

I meant the whole of